If a 0.0045 m^3 container has 1.5 moles of oxygen and is at a temperature of 303 K, what is its pressure in atm? (looks like '5.26 atm')

(Some previous posters were given an answer of 8.28, but that is not correct and has not been fixed on previous posts)
Answer #1

Per ideal gas equation,

PV=nRT

Let pressure be P

V=0.0045 m³

n=1.5

R=8.314

T=303 K

P=nRT/V=1.5*8.314*303/0.0045=839714 Pa=8.39 atm